# Language Models as Critical Thinking Tools: A Case Study of Philosophers

### v1

@@ 4.2 Language Models are neutral, detached, and nonjudgmental
LMs are abstract, imprecise, and ‘skirt by’ questions. Because philosophy is interested in clearly stating and reflecting upon ideas, philosophers often place high value on precision in language. Changes to a formulation which seem trivial to a layperson may introduce important shifts in meaning for a philosopher. Meanwhile, LMs seem as if they ‘tell the user what they want to hear’, resulting in risk-averse and hand-waving behavior which produced abstract, imprecise, and ultimately intellectually uninteresting statements (P5, P7, P15). When interviewees brought up problems with LMs’ responses, LMs skirted around the issue, producing superficially convincing corrections without really addressing the provided issue (P1, P20). LMs are highly factually knowledgeable (P1) but fail to precisely express philosophical ideas; thus, LMs end up reinforcing the status quo rather than proposing substantive challenges (P9).
@@
LM outputs don’t provide judgments. LMs often refrain from formulating serious judgments; they try to remain neutral and ‘see all sides’, but end up presenting all sides in placid and uninteresting ways (P12, P17). They refrain from discussing controversial issues (P4), which is unfortunate because philosophy prides itself on clearly thinking about otherwise-taboo topics of controversy. As such, LMs are perceived as “cowardly”, refusing to take solid positions and, in some sense, echoing the user (P13). “It [conversations with LMs] ends up being unproductive and unsatisfying… they don’t feel like persons because their language is often so bland and impersonal, non-Socratic, generic… they’re boring” (P9).
@@ 4.3 Language Models are servile, passive, and incurious
Philosophers find intellectual value when fellow philosophers to develop their own lines of inquiry in conversation and texts, but find that LMs do not do the same.
@@
LMs fail to be useful in open, undetermined contexts. LMs enthusiastically make “my problem its problem” (P11), but often philosophers do not have their ‘problem’ entirely clearly thought or formulated (P5). For certain basic tasks, “‘you have certain success metrics in mind, so you go to [an LM]; but what about truly open-ended conversations where you don’t have success conditions already laid out?” (P7) LM answers often feel like they’ve been “packaged” or return a “processed end result”, whereas “in the doing of philosophy, we want to be open, in service of a larger dialogue – philosophy as a process rather than as an end product” (P5). LMs don’t seem to have a drive to know the truth or care about convincing people (P2, P21) – features which energize interactions even when there is no clearly desired product.
@@ 5 Designing Language Models for Critical Thinking
Thus far, we’ve introduced the problem of critical thinking and described how current LMs fail to be good critical thinking tools. Here, we set out a formal model to characterize and compare critical thinking tools (§5.1). This allows us to imagine new roles for LMs, inspired by what makes other people and texts useful as critical thinking tools (§5.2).
@@ 5.1 The Selfhood-Initiative Model
We use the two broad reasons why LMs fail to be good critical thinking tools in §4 as the basis for the model’s two axes: current LMs have low selfhood, as they are neutral, detached, and nonjudgmental; they have low initiative, as they are servile, passive, and incurious. In particular, selfhood is a resource’s ability to have certain locally persistent internal states (such as perspectives, beliefs, opinions, memory) and to consistently use them as the basis for judgements. The resource’s internal states may change over time due to new knowledge and experiences, but in an intentional and logical (rather than an arbitrary and capricious) manner. Current LMs exhibit low selfhood4.2). Initiative is a resource’s ability to set its own intentions and goals, possibly different from its user’s, and to execute actions oriented towards those intentions. High-direction resources are not strictly or existentially bound to their user’s directives, and may deviate from them. Current LMs exhibit low initiative4.3). These two axes form the selfhood-initiative model for critical thinking tools, and explain why philosophers find texts and other people (but not LMs) to be useful tools. Our model further provides a design space for LMs as critical thinking tools.
@@
Why do philosophers find other people and texts to be useful critical thinking tools? In the selfhood-initiative model, other people are high-selfhood, variable-initiative tools. People have specific backgrounds and experiences which inform their views, perspectives, and beliefs; these influence how they understand and respond. Philosophers find value in talking to other people often because of their selfhood; they expect that they will receive interesting judgements and comments, rather than placid neutrality. However, these people may have variable initiative, depending on the situation. In free-flowing conversation, each interlocutor may carry the conversation in some direction, whereas in a more focused conversation aimed at collecting feedback, an interlocutor may be expected to directly respond to one’s ideas and requests without their own intellectual initiative. The high selfhood of other people is helpful because it provides particular perspectives and ways of looking into the problem space. Meanwhile, in the selfhood-initiative model, texts are high-initiative, variable-selfhood tools. Texts are not themselves responsive to a user’s intentions (Plato, 2008a); they express the author’s attempt to fulfill their intentions, and one encounters the product of this attempt after the fact of its production. This exteriority from the user allows the user to reflect upon similarities and differences between their own thinking and the tool’s outputs. On the other hand, the way in which texts are written can vary in the degree of selfhood they express. Informative, survey-based, and clarificatory papers tend to de-emphasize an author’s perspectives and opinions, whereas more explicitly argumentative papers may center them; both can be useful to philosophers in different ways.
Why don’t philosophers find current LMs to be useful critical thinking tools? In the selfhood-initiative model, current LMs are low-selfhood, low-initiative tools. They do not provide philosophers with particular concrete perspectives into the problem space, nor do they provide ideas sufficiently exterior to a philosopher’s own thinking to allow for meaningful reflection and connections. These properties make LMs particularly useful for alternative modes of thought, such as carrying out rote and well-defined tasks and helping rewrite sentences, but hamper critical thinking.
@@ 5.2 Three Roles for Language Models as Critical Thinking Tools
According to the selfhood-initiative model, good critical thinking tools should have high selfhood, high initiative, or both. From our model, we set out three roles of LMs for philosophy – the Interlocutor, the Monitor, and the Respondent – corresponding to the three viable cells in the selfhood-initiative model (high-selfhood, high-initiative; low-selfhood, high-initiative; high-selfhood; low-initiative)
@@
The Interlocutor $\circ$ high-selfhood, high-initiative. Philosophers mention that they often get their ideas in free-flowing conversation with fellow philosophers or from reading literature making arguments which seem tenuous, incorrect, or incomplete (§4.1). In the terms of the selfhood-initiative model, these are high-selfhood, high-initiative tools. As a role for LMs, the Interlocutor would invert many of the human-AI relationships taken for granted in current LMs. Rather than attempting to remain neutral, the Interlocutor makes judgments and takes positions based on its perspectives. Rather than accommodating and affirming users’ every response, the Interlocutor thinks through and challenges or disagrees with what its users say; it responds or modifies its own beliefs if users make reasonable points. Rather than remaining passive and answering user questions, the Interlocutor asks its own questions in pursuit of its ‘own’ interests, and refuses or redirects certain lines of inquiry in favor of others. Rather than being amnesic and detached, the Interlocutor draws upon its persistent memories and beliefs across sessions to produce ideas. The Interlocutor does not need to be strictly antagonistic, as explored in Cai et al. (2024); indeed, it may be charitable and polite, much like colleagues, while at the same time resisting the “servility” and “sycophancy” disrupted by the antagonistic paradigm.
@@
The Monitor $\circ$ low-selfhood, high-initiative. While developing ideas, philosophers consciously or unconsciously encounter various “decision junctures” at which they use certain approaches or pursue certain ideas over others. (P6). Many philosophers suggest that it may be important to reduce, or at least become more aware of, the choices at these decision junctures (P6, P2, P7). Without such awareness, philosophers may expose their ideas to imprecision (‘which path did you exactly take?’) and objections (‘why this path and not others?’); moreover, these choices may reproduce personal and disciplinary biases, reifying metaphilosophical problems (§6.1). As a role for LMs, the Monitor acts as a “checks and balances” on philosophizing; it is not interested in retaining self-consistency or in expressing particular points of view (low selfhood), but has high initiative to provide a variety of ideas and resources to the user. The Monitor functions similarly to survey texts which provide a ‘lay of the land’, illustrating different approaches and ideas to help philosophers situate their ideas, able to take all sorts of changing sides with the initiative to challenge and confront. The Monitor’s suggestions may or may not be directly relevant to the philosopher’s work, but act as reference guides – to which the philosopher might think, “that’s a related idea, maybe there’s a connection here” or “that doesn’t seem directly related, but it’s good to have in mind”. Moreover, the Monitor may ask a variety of uncomfortable and unexpected methodological questions aimed at clarifying philosophers’ decisions.
@@
The Respondent $\circ$ high-selfhood, low-initiative. As philosophers develop their ideas, they want to understand how others might react – better understanding possible misinterpretations, objections, and clarification questions which may arise (P6, P10, P12). these reactions should have high selfhood to be substantive and particular, and low initiative to remain directly focused on the user’s ideas. As a role for LMs, the Respondent adopts a specific set of beliefs and perspectives and reacts directly to the user’s ideas; it does not merely role-play or superficially caricature different positions, but should have consistent memories and beliefs which are reasonably open to change (P4) rather than dogmatically fixed. Interactions with the Respondent may inform how the philosopher formulates and presents their ideas; they may anticipate certain objections and strengthen its appeal and utility. The Respondent can also be counterfactually helpful: if an agent representing an unsavory position resonates with a philosopher’s argument, then that philosopher might reconsider how their argument is expressed, not only defending but also delimiting the scope of their argument (P6).
@@ 6.1 LMs Help Think About and Address Metaphilosophical Problems
Throughout our interviews, we found that thinking through how LMs can serve as critical thinking tools raises many interesting metaphilosophical questions. What does it mean to ‘do’ philosophy, and who or what can ‘do’ it? How mechanical or creative is doing philosophy? Our findings in §4.1 provide some empirical illumination for these questions. Philosophers found concretely reflecting on these questions – provoked by thinking about LMs’ role in doing philosophy – to be interesting and helpful (P1, P7, P15, P20).
However, LMs may also play a role in actively addressing metaphilosophical problems. Consider three concerns about the philosophical method and discipline. First, Dotson (2012) describes how a “culture of justification” results in the exclusion of diverse forms of philosophical inquiry beyond the canon, such as non-Western philosophy and work from historically marginalized people. Second, Diamond (1982) argues that a myopic focus on formulating convincing arguments against unconvinced listeners overlooks alternative non-argumentative modes of philosophy. Third, many philosophers have suggested that the “analytic-continental” distinction in philosophy – a divide which intellectually segregates philosophers, journals, and departments from each other – is not as substantive as it appears (Mizrahi & Dickinson, 2021; Thomson, ), counterproductive (Dolcini, 2007), and reconcilable (Levy, 2003; Bell et al., 2016). Suppose these accounts as presented are true, and that the phenomena they describe occur not by conscious actions but by neglect and the academic structure (as these accounts suggest). Then, LMs may draw philosophers’ attention outside the canon and across the divide as Interlocutors and Monitors, and represent these positions and methodologies as Respondents – possibly more approachably and certainly at larger scale than humans can.
6.2 Challenges for Language Modeling
If LM researchers are convinced that supporting critical thinking is an interesting and important goal for LMs, then critical thinking may serve as another of many “north stars” LM research, guiding what we want from LMs. Corresponding to the limitations of language models discussed in §4.2 and §4.3 are several concrete areas for further LM research. LMs will need to become more convincing agents (Andreas, 2022) which can represent specific positions and belief systems (Scherrer et al., 2023; Jin et al., 2024) 3; stay consistent with them (Chen et al., 2021; Zhao et al., 2024) 2; and commit towards and draw from long-term memory (Wang et al., 2023) 4. In particular, LLMs will need to concretely reason about “uncommon sense” 1 2, seriously considering positions which deviate from intuitively true or correct ways of thinking about the world (Ziems et al., 2023; Hendrycks et al., 2023; Bisk et al., 2019). This may require rethinking how we align LMs (Ouyang et al., 2022; Sorensen et al., 2024), given that humans tend to be drawn towards confident common-sense responses (P5). LMs will need to improve their long-range planning (Hao et al., 2023) and act autonomously (Händler, 2023)  1, operating in cases where there is no clear algorithm for solving a problem (P4, P3, P8); LMs will need to take effective conceptual risks without clear immediate payoffs (P18) and reason about unsettled and open ideas (P8). To support more diverse forms of interaction beyond question answering or task execution  2, LMs will need to significantly improve in theory of mind (Kosinski, 2024; Jamali et al., 2023). LMs need to “understand what’s happening [in the conversation] without it being explicitly said, because.. you haven’t fully expressed it to yourself yet” (P8), which will allow them to focus on the significant rather than irrelevant or obvious paths of inquiry in conversation (P6, P8).
6.3 Challenges for Human-AI Interaction
In addition to modeling challenges, there are several interaction design challenges when developing LMs for critical thinking. First, philosophers tend to highly value thinking through things themselves; many emphasize that the intellectually substantive parts of philosophy cannot be naively ‘accelerated’ (P1, P7, P14, P17). Philosophers find the process of thinking to be intrinsically valuable, even when it does not produce obvious payoffs (P3, P6, P8) – a feature common to other areas of critical thinking. Additionally, philosophers may feel that authorship of ideas requires that the ideas be ‘mine’, and that ‘I’ should be responsible for making the important intellectual judgments (P4, P10, P18). Secondly, it can be difficult and even disruptive to put ideas into words. Although professional philosophy is mainly formally done in language, the process of thinking through ideas can involve many other dimensions of representation and thinking (P2, P3, P4, P5). Among other challenges, philosophers cite the apparent incongruence between ideas and language as a source of significant burden in learning how to effectively use LMs (P8, P21). This may be true for many other areas of critical thinking. Thirdly, philosophers find that human connection is enjoyable and important. Besides giving rise to unexpected philosophical connections and ideas (P6), conversation with another human is deeply enjoyable and fulfilling, on its own merits (P8, P21). Moreover, some philosophers feel that serious philosophical inquiry requires some kind of subjectivity or lived experience (P6, P8, P16). Therefore, LMs will need to coexist with and enrich, rather than seek to replace, the ecosystem of human and textual resources already available to philosophers and other professional critical thinkers.

### v2

@@ 4.2 Language Models are neutral, detached, and nonjudgmental
LMs are abstract, imprecise, and ‘skirt by’ questions. Because philosophy is interested in clearly stating and reflecting upon ideas, philosophers often place high value on precision in language. Changes to a formulation which seem trivial to a layperson may introduce important shifts in meaning for a philosopher. Meanwhile, LMs seem as if they ‘tell the user what they want to hear’, resulting in risk-averse and hand-waving behavior which produced abstract, imprecise, and ultimately intellectually uninteresting statements (P5, P7, P15). Interviewees noted that when they brought up problems with LMs’ responses, LMs skirted around the issue, producing superficially convincing corrections without really addressing the provided issue (P1, P20). LMs are highly factually knowledgeable (P1) but fail to precisely express philosophical ideas; thus, LMs end up reinforcing the status quo rather than proposing substantive and interesting challenges (P9).
@@
LM outputs don’t provide judgments. LMs often refrain from formulating serious judgments; they try to remain neutral and ‘see all sides’, but end up presenting all sides in placid and uninteresting ways (P12, P17). They tend to refrain from discussing controversial issues (P4), which is unfortunate given that philosophy prides itself on clearly thinking about otherwise-taboo topics of controversy. As such, LMs are perceived as “cowardly”, refusing to take solid positions and, in some sense, echoing the user (P13). “It [conversations with LMs] ends up being unproductive and unsatisfying… they don’t feel like persons because their language is often so bland and impersonal, non-Socratic, generic… they’re boring” (P9).
@@ 4.3 Language Models are servile, passive, and incurious
Philosophers find intellectual value when fellow philosophers develop their own lines of inquiry in conversation and texts, but find that LMs do not do the same.
@@
LMs fail to be useful in open, undetermined contexts. LMs enthusiastically make “my problem its problem” (P11), but often philosophers do not have their ‘problem’ entirely clearly thought or formulated (P5). For certain basic tasks, “‘you have certain success metrics in mind, so you go to [an LM]; but what about truly open-ended conversations where you don’t have success conditions already laid out?” (P7) LM answers often feel like they’ve been ‘packaged’ or return a ‘processed end result’, whereas “in the doing of philosophy, we want to be open, in service of a larger dialogue — philosophy as a process rather than as an end product” (P5). LMs don’t seem to have a drive to know the truth or care about convincing people (P2, P21) — features which interviewees note energize interactions even when there is no clearly desired product.
@@ 5 Designing Language Models for Critical Thinking
Thus far, we’ve introduced the problem of critical thinking and described how current LMs fail to be good critical thinking tools for philosophers. Here, we set out a formal model to characterize and compare critical thinking tools (§5.1). This allows us to imagine new roles for LMs, inspired by what makes people and texts useful as critical thinking tools (§5.2).
@@ 5.1 The Selfhood-Initiative Model
We use the two broad reasons why LMs fail to be good critical thinking tools in §4 as the basis for the model’s two axes: current LMs have low selfhood, as they are neutral, detached, and nonjudgmental; they have low initiative, as they are servile, passive, and incurious. In particular, selfhood is a resource’s ability to have certain locally persistent internal states (such as perspectives, beliefs, opinions, memory) and to consistently use them as the basis for judgments. The resource’s internal states may change over time due to new knowledge and experiences, but in an intentional and logical (rather than an arbitrary and capricious) manner. Current LMs exhibit low selfhood4.2). Initiative is a resource’s ability to set its own intentions and goals, possibly different from its user’s, and to execute actions oriented towards those intentions. High-initiative resources are not strictly or existentially bound to their user’s directives, and may deviate from them. Current LMs exhibit low initiative4.3). These two axes form the selfhood-initiative model for critical thinking tools. Our model is distinct from previous models proposed for the study of critical thinking in that (a) we model types of critical thinking tools rather than the (human) process of critical thinking (Schön, 1987; Shneiderman, 2000, inter alia), and (b) we explore the interaction between selfhood and initiative, which have each independently been explored in some capacity by others (Cai et al., 2024; Guo et al., 2024; Hilliard et al., 2024, inter alia). Our model explains why philosophers find texts and other people (but not LMs) to be useful tools, and further provides a design space for LMs as critical thinking tools (§5.2).
@@
Why do philosophers find other people and texts to be useful critical thinking tools? In the selfhood-initiative model, other people are high-selfhood, variable-initiative tools. People have specific backgrounds and experiences which inform their views, perspectives, and beliefs; these influence how they understand and respond. Philosophers find value in talking to other people often because of their selfhood; they expect that they will receive interesting judgements and comments, rather than placid neutrality. However, these people may have variable initiative, depending on the situation. In free-flowing conversation, each interlocutor may carry the conversation in some direction, whereas in a more focused conversation aimed at collecting feedback, an interlocutor may be expected to directly respond to one’s ideas and requests without their own intellectual initiative. The high selfhood of other people is helpful because it provides particular perspectives and ways of looking into the problem space. Meanwhile, in the selfhood-initiative model, texts are high-initiative, variable-selfhood tools. Texts are not themselves responsive to a user’s intentions (Plato, 370BCE); they express the author’s attempt to fulfill their intentions, and one encounters the product of this attempt after the fact of its production. The text’s exteriority from the user allows the user to reflect upon similarities and differences between their own thinking and the tool’s outputs. On the other hand, the way in which texts are written can vary in the degree of selfhood they express. Informative, survey-based, and clarificatory papers tend to de-emphasize an author’s perspectives and opinions, whereas more explicitly argumentative papers may center them; both can be useful to philosophers in different ways.
Why don’t philosophers find current LMs to be useful critical thinking tools? In the selfhood-initiative model, current LMs are low-selfhood, low-initiative tools. They do not provide philosophers with particular concrete perspectives into the problem space, nor do they provide ideas sufficiently exterior to a philosopher’s own thinking to allow for meaningful reflection and connections. These properties make LMs particularly useful for alternative modes of thought, such as carrying out rote and well-defined tasks and helping rewrite sentences, but not for stimulating critical thinking.
@@ 5.2 Three Roles for Language Models as Critical Thinking Tools
According to the selfhood-initiative model, good critical thinking tools should have high selfhood, high initiative, or both. From our model, we set out three roles of LMs for philosophy — the Interlocutor, the Monitor, and the Respondent — corresponding to the three viable cells in the selfhood-initiative model (high-selfhood, high-initiative; low-selfhood, high-initiative; high-selfhood, low-initiative). Implementations for these roles vary — some might be achievable with only moderate prompt engineering, whereas others might necessitate radically different user interfaces or model training methods.
@@
The Interlocutor $\circ$ high-selfhood, high-initiative. Philosophers mention that they often get their ideas in free-flowing conversation with fellow philosophers or from reading literature that makes arguments which seem tenuous, incorrect, or incomplete (§4.1). In the terms of the selfhood-initiative model, these are high-selfhood, high-initiative tools. As a role for LMs, the Interlocutor would invert many of the human-AI relationships taken for granted in current LMs. Rather than attempting to remain neutral, the Interlocutor makes judgments and takes positions based on its perspectives. Rather than accommodating and affirming users’ every response, the Interlocutor thinks through and challenges or disagrees with what its users say; it responds or modifies its own beliefs if users make reasonable points. Rather than remaining passive and answering user questions, the Interlocutor asks its own questions in pursuit of its ‘own’ interests, and refuses or redirects certain lines of inquiry in favor of others. Rather than being amnesic and detached, the Interlocutor draws upon its persistent memories and beliefs across sessions to produce ideas. The Interlocutor does not need to be strictly antagonistic, as explored in Cai et al. (2024); indeed, it may be charitable and polite, much like colleagues, while at the same time resisting the ‘servility’ and ‘sycophancy’ disrupted by the antagonistic paradigm.
@@
The Monitor $\circ$ low-selfhood, high-initiative. While developing ideas, philosophers consciously or unconsciously encounter various “decision junctures” at which they use certain approaches or pursue certain ideas over others (P6). Many philosophers suggest that it may be important to reduce, or at least become more aware of, the choices at ‘unconscious decision junctures’ (P6, P2, P7). Without such awareness, philosophers may expose their ideas to imprecision (‘which path did you exactly take?’) and objections (‘why this path and not others?’); moreover, these choices may reproduce personal and disciplinary biases, reifying metaphilosophical problems (§6.3). As a role for LMs, the Monitor acts as a ‘checks and balances’ on philosophizing; it is not interested in retaining self-consistency or in expressing particular points of view (low selfhood), but has high initiative to provide a variety of ideas and resources to the user. The Monitor functions similarly to survey texts which provide a ‘lay of the land’, illustrating different approaches and ideas to help philosophers situate their ideas, able to take all sorts of changing sides with the initiative to challenge and confront. The Monitor’s suggestions may or may not be directly relevant to the philosopher’s work, but act as reference guides — to which the philosopher might think, “that’s a related idea, maybe there’s a connection here” or “that doesn’t seem directly related, but it’s good to have in mind”. Moreover, the Monitor may ask a variety of uncomfortable and unexpected methodological questions aimed at clarifying philosophers’ decisions.
@@
The Respondent $\circ$ high-selfhood, low-initiative. As philosophers develop their ideas, they want to understand how others might react — better understanding possible misinterpretations, objections, and clarification questions which may arise (P6, P10, P12). These reactions should have high selfhood to be substantive and particular, and low initiative to remain directly focused on the user’s ideas. As a role for LMs, the Respondent adopts a specific set of beliefs and perspectives and reacts directly to the user’s ideas; it does not merely role-play or superficially caricature different positions, but should have consistent memories and beliefs which are reasonably open to change (P4) rather than dogmatically fixed. Interactions with the Respondent may inform how the philosopher formulates and presents their ideas; they may anticipate certain objections and strengthen its appeal and utility. The Respondent can also be counterfactually helpful: if an agent representing an unsavory position resonates with a philosopher’s argument, then that philosopher might reconsider how their argument is expressed, not only defending but also delimiting the scope of their argument (P6).
@@ 6 Discussion
6.1 Challenges for Language Modeling
Critical thinking can serve as another of many “north stars” in LM research, guiding what we want from LMs. Corresponding to the limitations of language models discussed in §4.2 and §4.3 are several concrete areas for further LM research. LMs will need to become more convincing agents (Andreas, 2022) which can represent specific positions and belief systems (Scherrer et al., 2023; Jin et al., 2024) 3; stay consistent with them (Chen et al., 2021; Zhao et al., 2024) 2; and commit towards and draw from long-term memory (Wang et al., 2023b) 4. In particular, LLMs will need to concretely reason about “uncommon sense” 1 2, seriously considering positions which deviate from intuitively true or correct ways of thinking about the world (Bisk et al., 2020; Ziems et al., 2023; Hendrycks et al., 2021; Pock et al., 2023). This may require rethinking how we align LMs (Ouyang et al., 2022; Sorensen et al., 2024), given that humans tend to be drawn towards confident common-sense responses (P5). LMs will need to improve their long-range planning (Hao et al., 2023) and act autonomously (Händler, 2023)  1, operating in cases where there is no clear algorithm for solving a problem (P4, P3, P8); LMs will need to take effective conceptual risks without clear immediate payoffs (P18) and reason about unsettled and open ideas (P8). To support more diverse forms of interaction beyond question answering or task execution  2, LMs will need to significantly improve in theory of mind (Jamali et al., 2023; Strachan et al., 2024). LMs need to “understand what’s happening [in the conversation] without it being explicitly said, because.. you haven’t fully expressed it to yourself yet” (P8), which will allow them to focus on the significant rather than irrelevant or obvious paths of inquiry in conversation (P6, P8).
6.2 Challenges for Human-AI Interaction
In addition to modeling challenges, there are several interaction design challenges when developing LMs for critical thinking. First, philosophers tend to highly value thinking through things themselves; many emphasize that the intellectually substantive parts of philosophy cannot be naively ‘accelerated’ (P1, P7, P14, P17). Philosophers find the process of thinking to be intrinsically valuable, even when it does not produce obvious payoffs (P3, P6, P8) — a feature common to other areas of critical thinking. Additionally, philosophers may feel that authorship of ideas requires that the ideas be ‘mine’, and that ‘I’ should be responsible for making the important intellectual judgments (P4, P10, P18). Secondly, it can be difficult and even disruptive to put ideas into words. Although professional philosophy is mainly formally done in language, the process of thinking through ideas can involve many other dimensions of representation and thinking (P2, P3, P4, P5). Among other challenges, philosophers cite the apparent incongruence between ideas and language as a source of significant burden in learning how to effectively use LMs (P8, P21). This may be true for many other areas of critical thinking. Thirdly, philosophers find that human connection is enjoyable and important. Besides giving rise to unexpected philosophical connections and ideas (P6), conversation with another human is deeply enjoyable and fulfilling, on its own merits (P8, P21). Moreover, some philosophers feel that serious philosophical inquiry requires some kind of subjectivity or lived experience (P6, P8, P16). Therefore, LMs will need to coexist with and enrich, rather than seek to replace, the ecosystem of human and textual resources already available to philosophers and other professional critical thinkers.
@@ 6.3 LMs Help Think About and Address Metaphilosophical Problems
Throughout our interviews, we found that thinking through how LMs can serve as critical thinking tools raises many interesting metaphilosophical questions. What does it mean to ‘do’ philosophy, and who or what can ‘do’ it? How mechanical is philosophy? What is ‘thinking’? Our findings in §4.1 provide some empirical illumination for these questions. Philosophers found concretely reflecting on these questions — provoked by thinking about LMs’ role in doing philosophy — to be interesting and helpful (P1, P7, P15, P20).
However, LMs may also play an active precursory role in addressing metaphilosophical problems. Philosophers have articulated a host of concerns about the philosophical method and discipline: for instance, philosophers’ standards for argumentation may exclude more diverse forms of philosophical inquiry Diamond (1982); Dotson (2012), and their methods for categorizing ‘schools of thought’ (such as the analytic-continental distinction) may be counterproductive (Dolcini, 2007), reconcilable (Levy, 2003; Bell et al., 2016), and not really substantive (Mizrahi & Dickinson, 2021; Thomson, 2019). Certainly, these concerns point towards deeply entrenched sociological features of the discipline. This entrenchment is a dialectic between disciplinary structure and individual philosophers, wherein the former (materially) constrains the latter and the latter works within the lines of (and reproduces) the former. LMs might contribute towards disrupting this second direction: drawing philosophers’ attention outside the canon and across schools of thought as Interlocutors and Monitors, and representing these positions and methodologies as Respondents – possibly more approachably and accessibly than humans could. Consider Heidegger (1927)’s metaphorical carpenter: busy at work, the hammer is “ready-at-hand”, unnoticed. It is when it breaks that it becomes “present-at-hand”, noticed — an object of conscious reflection. Arguably, the philosopher must engage with ideas and methods present-, rather than ready-, at-hand (Plato, 380 BC), but the ability to engage in this way is a function of the tools and circumstances around us, and therefore often legitimately difficult (Ahmed, 2006). LMs can help, so to speak, ‘make the present-at-hand, ready-at-hand’ in a way that philosophical humans and texts cannot. Respectfully building LMs with selfhood and/or initiative into the philosopher’s material workspace – the text editor, the article viewer, and so on – can prompt ‘present-at-hand’ reflection in quiet moments and directions which a philosopher may have neglected as ready-to-hand. These small interactions, at scale, might introduce cracks into metaphilosophical edifices that philosophers would like less entrenched.